# AnesSuite: A Comprehensive Benchmark and Dataset Suite for Anesthesiology Reasoning in LLMs

### v2

@@ 1 Introduction
Leveraging AnesSuite, we propose Morpheus, the first collection of reasoning LLMs specifically designed for anesthesiology. Initialized from the Qwen2.5-7B and Qwen2.5-14B (Yang et al., 2024), it was trained on AnesR1 using SFT and GRPO (Shao et al., 2024). Despite undergoing limited training, experimental results demonstrate that Morpheus achieves substantial performance improvements on AnesBench, rivaling the performance of larger-scale models.
@@
Our contributions are threefold: (1) Dataset Suite Construction: We introduce AnesSuite, the first comprehensive dataset suite for anesthesiology reasoning. It includes AnesBench, a structured, cross-lingual benchmark for evaluating factual retrieval and clinical decision-making, along with three training datasets supporting CPT, SFT, and RLVR. (2) Baseline Model: We present Morpheus, the first collection of reasoning LLMs for anesthesiology. With SFT and GRPO training on the AnesR1 dataset, it achieves substantial performance improvements, underscoring the value of specialized, domain-specific training. (3) Comprehensive Evaluation and Analysis: Through systematic evaluations and extensive ablation studies, we provide an in-depth analysis of factors influencing anesthesiology reasoning, such as model characteristics, training strategies and training data. Overall, we are committed to leveraging our dataset suite, baseline models, and in-depth analysis collectively provide valuable insights for developing LLMs with more enhanced anesthesiology reasoning capabilities. We will publicly release AnesSuite, Morpheus, and all associated resources.
@@ 2.2 Anesthesia-related Benchmarks and LLMs
Anesthesiology, due to its unique interdisciplinary nature, is often implicitly classified under surgery, dentistry, or similar categories (Panagoulias et al., 2024; Olatunji et al., 2024; Abacha et al., 2019), even in benchmarks that categorize based on different departments. Some benchmarks explicitly treat anesthesiology as a separate category (Grzybowski et al., 2024; Wang et al., 2024a), while some studies employ existing anesthesiology-related questions (Khan et al., 2024; Angel et al., 2022; Shay et al., 2023; Kinoshita et al., 2024; Patnaik & Hoffmann, 2023) (e.g., exam questions from American Society of Anesthesiologists, ASA and Japanese Society of Anesthesiologists, JSA) to evaluate LLMs. However, the overall scale of these evaluations for anesthesiology remains very limited. Chinese Anesthesiology Benchmark (Zhou et al., 2025, CAB) is the first to focus primarily on anesthesiology. Nevertheless, CAB’s primary emphasis on factual retrieval questions results in insufficient exploration of the unique challenges in anesthetic reasoning and decision-making. Additionally, its restriction to the Chinese language further limits its applicability and impact. Moreover, Hypnos and some works (Wang et al., 2024b; Guthrie et al., 2024) attempts to use anesthesiology-related corpora for supervised fine-tuning to improve question-answering abilities, but they have not significantly enhanced the reasoning capabilities.
@@ 4.2 Baseline Model
We developed baseline models by fine-tuning Qwen2.5-7B and Qwen2.5-14B (Yang et al., 2024) on the AnesR1. Both models underwent SFT followed by GRPO and are designated Morpheus-7B and Morpheus-14B, respectively. For comparison, we also present counterpart models trained for an equivalent number of steps using SFT alone. Complete training details are provided in Appendix C.2.
The evaluation results for Morpheus are presented in Tab. 4. A key observation is that, after GRPO training, the model achieves consistent improvements across various systems, demonstrating the effectiveness of the GRPO algorithm in enhancing reasoning capabilities. Notably, following SFT and GRPO training on AnesR1, Morpheus-7B and Morpheus-14B reach performance levels comparable to Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct, respectively. This further indicates that AnesR1 is effective in advancing complex anesthesiology reasoning abilities.
@@ 7 Conclusion
This paper introduces AnesSuite, the first comprehensive suite dedicated to anesthesiology reasoning. AnesSuite features a structured evaluation benchmark and three training datasets, establishing a complete infrastructure for key development stages: CPT, SFT, and RLVR. We also developed Morpheus, the first strong baseline model collection for this domain. Furthermore, extensive ablation studies and analyses yield several key insights for improving these specialized reasoning capabilities. We hope this work will serve as a cornerstone for developing LLMs with enhanced anesthesiology reasoning capabilities.
@@ C.2 Experiment Details for Baseline Model
In the second stage, we employed the Verl framework (Sheng et al., 2025) for model training using GRPO. The primary training parameters were configured as follows: a training batch size of 512 and a mini-batch size of 256 for each gradient update. The maximum response length was constrained to 2048 tokens. For each prompt, the group size for rollout was set to 5. The training process was conducted for a maximum of 4 epochs, with a learning rate of $1\times 10^{-6}$ and a learning rate warmup ratio of 0.3. Under this configuration, the Morpheus-7B and Morpheus-14B models were trained for 50 and 40 steps, respectively.
@@ Appendix G Supplementary Evaluation
G.1 Self-Consistency in Test Time Scaling
@@
G.2 Evaluation on the Open-ended Subset

### v3

@@ 1 Introduction
Leveraging AnesSuite, we propose Morpheus, the first collection of reasoning LLMs specifically designed for anesthesiology. Initialized from the Qwen2.5-7B, Qwen2.5-14B and Qwen2.5-32B (Yang et al., 2024), it was trained on AnesR1 using SFT and GRPO (Shao et al., 2024). Despite limited training, Morpheus consistently enhances reasoning performance across domain-specific, general medical, and general-domain benchmarks, rivaling the performance of larger-scale models.
@@
Our contributions are threefold: (1) Dataset Suite Construction: We introduce AnesSuite, the first comprehensive dataset suite for anesthesiology reasoning. It includes AnesBench, a structured, cross-lingual benchmark for evaluating factual retrieval and clinical decision-making, along with three training datasets supporting CPT, SFT, and RLVR. (2) Baseline Model: We present Morpheus, the first collection of reasoning LLMs for anesthesiology. With SFT and GRPO training on the AnesR1 dataset, Morpheus yields consistent reasoning enhancements across domain-specific, general medical, and general-domain benchmarks, underscoring the efficacy of data from reasoning-intensive disciplines like anesthesiology. (3) Comprehensive Evaluation and Analysis: Through systematic evaluations and extensive ablation studies, we provide an in-depth analysis of factors influencing anesthesiology reasoning, such as model characteristics, training strategies and training data. Overall, we are committed to leveraging our dataset suite, baseline models, and in-depth analysis collectively provide valuable insights for developing LLMs with more enhanced anesthesiology reasoning capabilities. We will publicly release AnesSuite, Morpheus, and all associated resources.
@@ 2.2 Anesthesia-related Benchmarks and LLMs
Anesthesiology, due to its unique interdisciplinary nature, is often implicitly classified under surgery, dentistry, or similar categories (Panagoulias et al., 2024; Olatunji et al., 2024; Abacha et al., 2019), even in benchmarks that categorize based on different departments. Some benchmarks explicitly treat anesthesiology as a separate category (Grzybowski et al., 2024; Wang et al., 2024a), while some studies employ existing anesthesiology-related questions (Khan et al., 2024; Angel et al., 2022; Shay et al., 2023; Kinoshita et al., 2024; Patnaik & Hoffmann, 2023) (e.g., exam questions from American Society of Anesthesiologists, ASA and Japanese Society of Anesthesiologists, JSA) to evaluate LLMs. However, the overall scale of these evaluations for anesthesiology remains very limited. Chinese Anesthesiology Benchmark (Zhou et al., 2025, CAB) is the first to focus primarily on anesthesiology. Nevertheless, CAB’s primary emphasis on factual retrieval questions results in insufficient exploration of the unique challenges in anesthetic reasoning and decision-making. Additionally, its restriction to the Chinese language further limits its applicability and impact. Moreover, Hypnos and some works (Wang et al., 2024c; Guthrie et al., 2024) attempts to use anesthesiology-related corpora for supervised fine-tuning to improve question-answering abilities, but they have not significantly enhanced the reasoning capabilities.
@@ 4.2 Baseline Model
We developed baseline models by fine-tuning Qwen2.5-7B, Qwen2.5-14B and Qwen2.5-32B (Yang et al., 2024) on the AnesR1. Both models underwent SFT followed by GRPO and are designated Morpheus-7B, Morpheus-14B and Morpheus-32B, respectively. In alignment with established paradigms, the SFT stage serves as a cold-start initializing for the subsequent GRPO training. For comparative purposes, we also present these intermediate models, which obtained after the limited SFT steps. Complete training details are provided in Appendix C.2. In addition, the results of AnesR1 on more models can be found in Tab. 11.
The evaluation results for Morpheus are presented in Tab. 4. A key observation is that, after GRPO training, the model achieves consistent improvements across various systems, demonstrating the effectiveness of the GRPO algorithm in enhancing reasoning capabilities. Notably, following SFT and GRPO training on AnesR1, Morpheus-7B, Morpheus-14B and Morpheus-32B reach performance levels comparable to Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct and Qwen2.5-72B-Instruct, respectively. This highlights the efficacy of AnesR1 in cultivating complex reasoning within the specialized domain of anesthesiology. Furthermore, as shown in Tab 17 (Appendix G.1), Morpheus exhibits enhanced reasoning capabilities across a wide range of general medical and general-domain benchmarks (e.g., MMLU, MedQA). Given that AnesR1 focuses exclusively on the niche field of anesthesiology, these results may offer new insights into the the field of reasoning generalization (Cheng et al., 2025).
@@ 7 Conclusion
This paper introduces AnesSuite, the first comprehensive suite dedicated to anesthesiology reasoning. AnesSuite features a structured evaluation benchmark and three training datasets, establishing a complete infrastructure for key development stages: CPT, SFT, and RLVR. We also developed Morpheus, the first strong baseline model collection for this domain. Trained exclusively on the AnesR1, these models demonstrate consistent performance gains across anesthesiology, general medical, and general-domain benchmarks. Furthermore, extensive ablation studies and analyses yield several key insights for improving these specialized reasoning capabilities. We hope this work will serve as a cornerstone for developing LLMs with enhanced anesthesiology reasoning capabilities.
@@ C.2 Experiment Details for Baseline Model
In the second stage, we employed the Verl framework (Sheng et al., 2025) for model training using GRPO. The primary training parameters were configured as follows: a training batch size of 512 and a mini-batch size of 256 for each gradient update. The maximum response length was constrained to 2048 tokens. For each prompt, the group size for rollout was set to 5. The training process was conducted for a maximum of 4 epochs, with a learning rate of $1\times 10^{-6}$ and a learning rate warmup ratio of 0.3. Under this configuration, the Morpheus-7B, Morpheus-32B and Morpheus-14B models were trained for 50, 50 and 40 steps, respectively. For the Morpheus-32B, due to the constraints of computing resources, we adopted the parameter-efficient fine-tuning method of LoRA. Among them, the rank $r$ of matrix factorization was set to 128, and the scaling factor $\alpha$ was set to 32.
In our specific training process, we exclusively employed an outcome-based binary reward mechanism: a reward of $1$ is assigned if and only if the model’s final output matches the ground truth; otherwise, the reward is 0. To accurately extract the final answer from the model’s output, we utilized robust regular expressions to strictly match patterns such as “Answer is” or “Answer:”. To demonstrate the stability of the training process, we conducted three independent training runs for the Morpheus-7B model. As illustrated by the reward curves in Fig 10 and Fig 10, the training process exhibits significant stability across both the 7B and 14B models.
Figure 9: Reward curve of Morpheus7B
Figure 10: Reward curve of Morpheus14B
Furthermore, we trained multiple models using the AnesR1 dataset with the number of training steps extended to 200. As shown in Tab 11, the results demonstrate significant performance improvements across both the Chinese and English subsets of AnesBench, validating the high quality and rich knowledge encapsulated within AnesR1. Notably, as shown in Tab. 4, the Qwen model exhibited a decline in Chinese subset following SFT, which may be attributed to its inherent sensitivities.
Model AnesBench-English AnesBench-Chinese Avg.
Llama-3.1-8B-Instruct 0.54 0.48 0.51
Llama-3.1-8B(AnesR1) 0.60 0.51 0.55
Gemma2-2B-Instruct 0.40 0.34 0.37
Gemma2-2B(AnesR1) 0.45 0.41 0.43
Gemma3-4B-Instruct 0.32 0.40 0.36
Gemma3-4B(AnesR1) 0.36 0.43 0.39
Table 11: Additional Training Experiments with AnesR1
@@ Appendix G Supplementary Evaluation
G.1 Evaluation on Out of Distribution Benchmarks
To further assess the generalization capability of the model trained on anesthesiology reasoning data, we evaluated Morpheus across three general medical benchmarks (MedQA, MedMCQA, and MedExpQA; Jin et al., 2021; Pal et al., 2022; Alonso et al., 2024) and three general-domain benchmarks (MMLU, MMLU-Pro, and AGIeval; Wang et al., 2023a; 2024b; Zhong et al., 2024). Consistent with our previous experiments, we employed the same framework used in AnesBench to extract answers for multiple-choice questions. Notably, for AGIeval, we restricted the evaluation to its 18 MCQ tasks, whereas comprehensive evaluations were conducted for all other benchmarks. The medical subset of MMLU includes the tasks anatomy, clinical_knowledge, college_biology, college_medicine, high_school_biology, medical_genetics, nutrition, professional_medicine, and virology.
Model AGIEval MMLU MMLU-Pro MedExpQA MedMCQA MedQA
Qwen2.5-7B-Instruct 0.649 0.709 0.455 0.526 0.532 0.654
Morpheus-7B 0.663 0.786 0.580 0.582 0.593 0.720
Qwen2.5-14B-Instruct 0.701 0.801 0.627 0.642 0.593 0.735
Morpheus-14B 0.747 0.843 0.667 0.752 0.661 0.798
Qwen2.5-32B-Instruct 0.739 0.828 0.672 0.746 0.628 0.780
Morpheus-32B 0.766 0.873 0.717 0.798 0.685 0.814
Qwen2.5-72B-Instruct 0.742 0.852 0.703 0.790 0.674 0.812
Table 17: Evaluation Result on OOD Benchmarks
Model English French Italian Spanish
Qwen2.5-7B-Instruct 0.600 0.512 0.512 0.480
Morpheus-7B 0.632 0.608 0.464 0.624
Qwen2.5-14B-Instruct 0.712 0.576 0.632 0.648
Morpheus-14B 0.736 0.728 0.776 0.768
Qwen2.5-32B-Instruct 0.784 0.768 0.712 0.720
Morpheus-32B 0.784 0.808 0.808 0.792
Qwen2.5-72B-Instruct 0.800 0.776 0.776 0.808
Table 18: Evaluation Result on MedExpQA
Model Mainland Taiwan US
Qwen2.5-7B-Instruct 0.797 0.660 0.505
Morpheus-7B 0.795 0.709 0.656
Qwen2.5-14B-Instruct 0.816 0.745 0.643
Morpheus-14B 0.835 0.815 0.746
Qwen2.5-32B-Instruct 0.853 0.804 0.682
Morpheus-32B 0.831 0.817 0.794
Qwen2.5-72B-Instruct 0.861 0.833 0.742
Table 19: Evaluation Result on MedQA
Model AQuA-RAT Gaokao JEC LSAT LogiQA SAT
Qwen2.5-7B-Instruct 0.622 0.740 0.563 0.476 0.528 0.722
Morpheus-7B 0.721 0.737 0.575 0.489 0.530 0.766
Qwen2.5-14B-Instruct 0.752 0.767 0.601 0.533 0.621 0.798
Morpheus-14B 0.772 0.817 0.670 0.585 0.648 0.830
Qwen2.5-32B-Instruct 0.744 0.817 0.663 0.576 0.652 0.804
Morpheus-32B 0.787 0.820 0.696 0.628 0.705 0.843
Qwen2.5-72B-Instruct 0.752 0.812 0.676 0.577 0.663 0.815
Table 20: Evaluation Result on AGIeval
Model medical College High_School Other Professional
Qwen2.5-7B-Instruct 0.728 0.595 0.773 0.703 0.581
Morpheus-7B 0.782 0.734 0.859 0.773 0.673
Qwen2.5-14B-Instruct 0.797 0.724 0.864 0.796 0.693
Morpheus-14B 0.842 0.791 0.905 0.834 0.741
Qwen2.5-32B-Instruct 0.827 0.801 0.896 0.812 0.732
Morpheus-32B 0.862 0.875 0.922 0.862 0.789
Qwen2.5-72B-Instruct 0.857 0.810 0.912 0.837 0.771
Table 21: Evaluation Result on MMLU
Task Q-7B-It M-7B Q-14B-It M-14B Q-32B-It M-32B Q-72B-It
Health 0.502 0.630 0.643 0.703 0.686 0.737 0.698
Biology 0.695 0.738 0.787 0.827 0.829 0.861 0.849
Business 0.459 0.649 0.710 0.693 0.734 0.771 0.769
Chemistry 0.413 0.572 0.657 0.627 0.691 0.730 0.736
Computer Sci 0.400 0.637 0.649 0.749 0.698 0.773 0.756
Economics 0.556 0.679 0.719 0.778 0.751 0.820 0.780
Engineering 0.366 0.339 0.490 0.396 0.550 0.474 0.542
History 0.417 0.533 0.541 0.625 0.583 0.640 0.619
Law 0.299 0.326 0.341 0.445 0.406 0.501 0.465
Math 0.486 0.759 0.778 0.810 0.799 0.869 0.826
Other 0.432 0.556 0.590 0.672 0.648 0.703 0.695
Philosophy 0.361 0.453 0.529 0.629 0.581 0.657 0.611
Physics 0.400 0.590 0.657 0.642 0.719 0.728 0.739
Psychology 0.589 0.654 0.691 0.734 0.727 0.773 0.756
Table 22: Evaluation Result on MMLU-Pro
The Morpheus-7B/14B/32B models, fine-tuned via SFT and GRPO on anesthesia reasoning data, consistently outperform the Qwen2.5-Instruct baselines across all three out-of-distribution general medical benchmarks. In certain settings, they even surpass the performance of 72B-scale models, indicating that our training strategy primarily enhances general medical reasoning capabilities rather than merely causing overfitting to the source task. Regarding language transferability, Morpheus generally outperforms the corresponding Qwen2.5-Instruct models across four languages on the MedExpQA benchmark. The improvements are particularly significant in Italian and Spanish, demonstrating strong cross-lingual reasoning transfer. A similar trend is observed in MedQA, where performance gains are notably substantial in Traditional Chinese and English contexts.
In terms of general-domain generalization, Morpheus demonstrates strong performance on benchmarks such as AGIEval, MMLU, and MMLU-Pro. The performance improvements observed across the majority of tasks indicate that the reasoning gains derived from our training data extend beyond the medical domain to general-purpose tasks.
Since we performed no training or fine-tuning on domains other than anesthesiology, no new domain-specific knowledge was introduced. Consequently, these performance gains can be attributed to enhanced reasoning capabilities, allowing the model to more effectively utilize its internal knowledge base. While prior works (Cheng et al., 2025) have focused on multi-domain reasoning generalization, our findings offer new insights into generalization derived from a highly specific, smaller source domain.
G.2 Self-Consistency in Test Time Scaling
@@
G.3 Evaluation on the Open-ended Subset